1. (4 points) A buffer solution is created with C2H5NH 4.7 x 104 and 0.012 M C2H5NH2. The Kb for C2H5NH2 1s -3 What is the coPPlease help me by checking part a, helping me do part b and c. Please show your work.
Answer #1

a)

Firstly, we will calculate pKb.

pk = -log K = -log (4.7 x 10-4) = 3.33

Then we will find pOH= 14- pH = 14-11.50 = 2.50

Then, we will use Henderson Hasselbalch equation,

pOH = pky + log- Salt (C2H5NH7] 1 = pk + log Base C2H5NH2

Plugging in the values, we have,

[C2H5NH3] 2.50 = 3.33+ log 0.012M [C2H5NH3] log 0.012M = -0.83 C2H5NH3] _ 10-0.83 0.012M C2H5NH 1 = 0.012M X 10-0.83 = 1.775

Thus, your calculation is correct.

b)

HCl will convert the base to its salt.

C2H5NH2 + H+ + C2H5NH,

Thus, [C2H5NH2] = 0.012 M - 0.0011 M = 0.0109 M

[C2H5NH3+]=0.0018 M + 0.0011 M= 0.0029 M

c)

The buffer is a basic buffer because the pH is above 7.0.

Yes, this buffer can create an acidic buffer as well depending on concentration of species. For Acidic buffer pH<7 or pOH>7. This will give,

POH > 7 (C2H5NH3) →pky + log [C2H5NH2] [CH NH] 3.33+ log [C2H5NH2] C2H5NH3) > 3.67 log C,H5NH2] C2H5NH3 103.67 [C2H5NH2] C2H5

If this condition is satisfied then the "buffer" will theoretically act as an acidic buffer.